OpenStudy (anonymous):

Which expression is equivalent to –(–2m + 1)? (Points : 1) –3m 3m –2m – 1 2m – 1

OpenStudy (anonymous):

It would be c

OpenStudy (texaschic101):

do you know how to distribute through the parenthesis ?

OpenStudy (anonymous):

because by the distributive property -(-2m) becomes (2m) and -(+1) becomes -1

OpenStudy (texaschic101):

not C

OpenStudy (anonymous):

Why isn't it c?

OpenStudy (anonymous):

Oh yeah

OpenStudy (anonymous):

It's D

OpenStudy (anonymous):

Thanks @texaschic101

OpenStudy (texaschic101):

- (- 2m + 1) = 2m - 1.....you have to distribute the - through the parenthesis

OpenStudy (anonymous):

What is the area of a circle whose radius is 20 meters? (Points : 1) 20 m2 40 m2 200 m2 400 m2 could you help with this one too

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

so you need to use the equation \[3.14r^2\]

OpenStudy (texaschic101):

A = pi * r^2 A = 3.14 * 20^2 you can drag out the calculator and do this

OpenStudy (anonymous):

Or\[3.14(20^2)\]

OpenStudy (anonymous):

So would it be A?

OpenStudy (anonymous):

No it would be d

OpenStudy (texaschic101):

it is actually pi * 400...there just not putting in pi

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Medal? @NashGriersGirly